# Survey on Modeling of Articulated Objects

### v1

@@ 3 Datasets
Recent advancements in deep learning have led to the creation of large-scale 3D datasets for various computer vision tasks, particularly in shape understanding. However, datasets for articulated objects are relatively scarce compared to those for static 3D objects. Gathering data for articulated objects is more laborious because it requires detailed modeling of each part’s geometry and careful annotation of articulation parameters. The challenges in creating datasets for articulated objects are multifaceted:
  • The creation of synthetic datasets involves manually designing the geometry of each component. This labor-intensive process restricts the number of categories and instances in these datasets, often leading to a lack of diversity and complexity in the data samples. As a result, models may be less realistic or overly simplistic.
  • Real-world datasets are typically gathered using sensors like RGB-D cameras, followed by post-processing to reconstruct shapes and annotate part attributes. One of the significant challenges in modeling articulated objects, as opposed to static surface models, is the difficulty in capturing the interior structure of each part from surface data alone. The data obtained from surface sensors is often incomplete and noisy, with the interior structure frequently occluded by the outer surface. This limitation impacts the geometric quality of the data and further reduces the diversity and complexity of data samples.
With the recent surge in interest in articulated objects, there has been a growing effort to create datasets for this domain. We collect the existing datasets for articulated objects in both object-level and scene-level in Table 1.
@@ 3.1 Synthetic Data
The recent advancements in 3D modeling technology have led to a significant increase in the availability of 3D models with part-level structures accessible online. This progress has opened up new possibilities for detailed analysis and application in various fields, enhancing the resources for researchers and practitioners working with complex 3D structures. Articulated objects are one such class of complex 3D structures that have benefited from this progress. The availability of 3D datasets with part-level structures has enabled the creation of synthetic datasets for articulated objects.
@@
To facilitate the prediction of part mobility via a learning-based approach, [HLV*17, ] collected a synthetic dataset for 3D articulated objects by sourcing shapes from ShapeNet [CFG*15] and SketchUp [Inc17a]. In this process, they manually segmented the shapes into parts and provided detailed annotations for the articulation parameters. Following this foundational work, RPM-Net [YHY*19] extends the scope of the dataset to include a greater variety of objects incorporating more complex motion, e.g. the opening of the umbrella cover. For each shape in the dataset, each pair of possible parts is labeled as a reference part and a moving part. And each pair is taken as a mobility unit, which is associated with annotated motion parameters. These parameters form a quadruple, including the transformation type (which can be translation, rotation, or the combination of the two), the position and direction of the transformation axis, and the range of the motion.
@@
Shape2Motion [WZS*19] is another synthetic dataset in a larger scale that was introduced concurrently. The shapes are sourced from ShapeNet [CFG*15] and 3D Warehouse [Inc17]. Each shape includes part segmentations and articulation parameters that are annotated in a similar way as RPM-Net [YHY*19] and [HLV*17, ]. However, the process is further refined with the use of a developed annotation tool. This tool enhances the efficiency of the annotation process by allowing the annotators to visually verify the correctness by animating the object with the annotated parameters. This allows for boosting the scale of the dataset with a much larger number of objects and movable parts in a wide range of categories.
@@
The primary application of the datasets mentioned earlier involves sampling point clouds from the mesh surfaces to analyze part mobility. However, a notable limitation of these datasets is the absence of texture information. This lack of texture detail restricts their utility in applications such as simulation, the creation of digital twins, augmented reality, and other modeling tasks that require a more holistic visual representation. To address this gap, the PartNet-Mobility dataset, with a simulation environment SAPIEN  [XQM*20], was introduced. This dataset is a subset and an extension from a part-level 3D shape dataset PartNet [MZC*19], enriched with articulation annotations organized in URDF files. The inclusion of diverse appearances and motions in PartNet-Mobility significantly enhances its value, inspiring increasingly more research in the field of articulated object modeling. Building upon the PartNet-Mobility dataset, OPDSynth [JMSC22] is introduced for the task of openable part detection (OPD) by blending the rendering of the synthetic data with real-world RGB images. This dataset enables the mobility part detection from RGB images, aligning more closely with practical real-world applications.
@@ 3.2 Real-world Data
A primary challenge in working with synthetic data is the often unrealistic assumption that it is perfect, overly simplistic, complete, and free from noise. To fill this synthetic-real gap, more recent efforts have started to focus on collecting data from real-world sensors, such as RGB-D cameras.
@@
RBO [MEB19] represents the pioneering dataset in this regard, being the first to collect data from RGB-D scans. It not only provides RGB-D recordings of human interaction with objects under varying experimental conditions, but also creates corresponding synthetic mesh models for each object. ReArt-48 [LXX*22] introduced a slightly larger dataset that reconstructs meshes from RGB-D scans and provides the articulation annotations across five object categories. Taking this further, AKB-48 [LXF*22] emerges as the first large-scale dataset for articulated objects based on real scans, where each object is described in a knowledge graph. To construct this dataset, a fast articulation knowledge modeling pipeline is presented, significantly reducing the cost and effort required for object modeling in the real world. This innovation enables the creation of 3D models on a scale comparable to PartNet-Mobility. Additionally, AKB-48 also annotates the physical properties of the objects, such as mass, to enhance the dataset’s applicability. This additional annotation is vital for bridging the generalization gap between simulation and real-world applications.
@@
In parallel, the MultiScan [MZJ*22] dataset marks a groundbreaking advancement as the first large-scale scene-level dataset that documents multiple states of articulated objects in indoor settings. To compile this dataset, a scalable 3D environment acquisition pipeline is designed for processing raw RGB-D scans to produce 3D surface mesh reconstruction with texture and articulation annotations for each articulated object in the scene. MultiScan is an invaluable resource for advancing research and applications that require an understanding of the dynamics and interactions of articulated objects in real-world environments at the scene level. Leveraging the MultiScan dataset, OPDMulti [SJSC23] is introduced for the OPD task of the multiple-parts version by extracting frames from the RGB-D scans and annotating the openable parts in each frame. It extends the OPD task to the scene level, which is more challenging and practical in real-life scenarios.
More recently, the GAPartNet [GXZ*23] dataset is introduced to capture finer-level part details that have been previously overlooked in existing datasets. This dataset is tailored to enhance the generalization of object perception and manipulation tasks across various object categories by focusing on the concept of Generalizable and Actionable Parts. The underlying premise is that functional parts, such as buttons and handles, are fundamental elements whose identification and extraction can significantly improve generalizability within and across object categories. They argue that the functional parts such as buttons and handles are more elementary, and the extraction of these parts can improve the generalizability in an intra-category manner. To achieve a comprehensive and versatile dataset, GAPartNet selectively compiles data from both the PartNet-Mobility [XQM*20] and AKB-48 [LXF*22] datasets to encompass both synthetic and real data. It provides a rich resource for robust algorithms that are applicable across scenarios in the realm of articulated object modeling and manipulation.
@@
Slightly different from the previous datasets, the OPDReal [JMSC22] dataset is introduced for the OPD task by collecting RGB-D scans from real-world scenes. By providing a set of real RGB images capturing various articulation states of the objects, along with reconstructed mesh models, OPDReal presents new opportunities for research in the field of articulated object modeling from 2D visuals. OPDMulti [SJSC23] follows this trend and extends the OPD task to multiple objects in the scene by collecting RGB-D frames, which is more challenging and practical in real-life scenarios.
@@
geo rep. # states part seg # parts partial/noisy aligned intermediate rep. strategy supervision geo rep. part seg
@@
[XWY*09, ] mesh 1 - surface patch handcrafted - mesh
[MYY*10, ] mesh 1 - surface patch handcrafted - mesh
[SHL*14, ] mesh multi surface patch handcrafted - mesh
[YLX*16, ] PC multi 3D trajectory handcrafted - PC
[LWL*16, ] RGB-D multi 3D trajectory handcrafted - PC
[HLV*17, ] mesh 1 - surface patch - - -
[YHL*18, ] PC 2 PC feat. supervised labeled PC PC
Shape2Motion [WZS*19] PC 1 - PC feat. supervised labeled PC PC
RPM-Net [YHY*19] PC 1 - PC feat. supervised labeled PC PC
[ATK19, ] RGB-D 1 - image feat. supervised labeled PC PC
[LWY*20, ] PC 1 - NAOCS, NPCS supervised labeled PC PC
[SCZ21, ] PC multi 3D trajectory - - PC
ScrewNet [JLCN21] PC multi PC feat. - - -
[AFA*22, ] PC 1 - PC feat. supervised labeled PC PC
[QJR*22, ] RGB multi image feat. supervised labeled PC 3D plane
OPD [JMSC22] RGB 1 - - image feat. supervised 2D mask 2D mask
OPDMulti [SJSC23] RGB 1 - - image feat. supervised 2D mask 2D mask
[LXX*22, ] RGB-D 1 - NPCS supervised labeled PC PC
GAPartNet [GXZ*23] RGB-D 1 - NPCS supervised labeled PC PC
[LZH*23, ] PC 1 - PC feat. self-supervised PC corr. PC
[LSH*23, ] PC 1 *** - PC feat. supervised labeled PC PC
Banana [DLS*23] PC 1 - PC feat. supervised labeled PC PC
@@
[PG08, ] PC multi *** PC handcrafted - mesh
A-SDF [MQK*21] SDF 1 neural field supervised SDF mesh
Ditto [JHZ22] PC 2 neural field supervised labeled PC mesh
CLA-NeRF [TLYS22] MV RGB 1 - neural field supervised 2D mask mesh
[WCM*22, ] MV RGB 1 - neural field supervised MV RGB mesh
CARTO [HIZ*23] MV RGB 1 - - neural field supervised SDF mesh
PARIS [LMS23] MV RGB 2 - neural field self-supervised MV RGB mesh
NAP [LDS*23] - - - - neural field generative - mesh
CAGE [LTMS23] - - - - bbox generative - mesh
Table 2: Summary of the related works in the aspect of geometry processing. The table provides information about the geometric representation and the number of articulted states observed in the input, the assumptions made on the input data, the methodology used in the geometry processing, and the output data representation. Please refer to section 4.2.1 and fig. 5 for the explanation of the intermediate representations.
4 Geometry Processing
In this section, we discuss the recent works on articulated object modeling from the perspective of geometry processing. Based on the goal of the tasks, we categorize the works into two groups: articulated part perception and articulated object creation. The former group of works focuses on understanding the articulated part structure of the objects from either 3D geometry or 2D observations. The latter group focuses on reconstructing or generating the 3D geometry of the articulated objects with either articulation parameters estimated from input observations or the ability to animate the objects to different articulation states. In Table 2, we summarize these works in terms of the choice of geometric representation, the assumptions that each work made for the input, and the methodology each work adopted for shape analysis. We specify the meaning of each column in Table 2 as follows:

### v3

@@ 3 Datasets
One of the main contributors to recent advances in deep learning is the availability of large-scale 3D datasets for various computer vision tasks, particularly in shape understanding. However, datasets for articulated objects are relatively scarce compared to those for static 3D objects. Gathering data for articulated objects is more laborious because it requires detailed modeling of each part’s geometry and careful annotation of articulation parameters. The challenges in creating datasets for articulated objects are multifaceted:
  • The creation of synthetic datasets involves manual design of the geometry for each component and annotating the articulation parameters. This labor-intensive process restricts the number of categories and instances in these datasets, often leading to a lack of diversity and complexity in the data samples. As a result, models may be less realistic or overly simplistic.
  • Real-world datasets are typically gathered using sensors like RGB-D cameras, followed by post-processing to reconstruct shapes and annotate part attributes. One of the significant challenges in modeling articulated objects, as opposed to static surface models, is the difficulty in capturing the interior structure of each part from surface data alone. It usually involves multiple scans from different states of the object to capture the complete geometry. The data obtained from surface sensors is often incomplete and noisy, as the interior structure frequently is occluded by the outer surface. This limitation affects the geometric quality of the data and further reduces the diversity and complexity of data samples.
With the recent surge in interest in articulated objects, there has been a growing effort to create datasets for this domain. Figure 4 shows several representative datasets for 3D articulated objects. We list all existing datasets for articulated objects in both object-level and scene-level in Table 1, where we summarize the key characteristics of each dataset. We also provide a brief overview of each dataset in the folllowing subsections.
@@ 3.1 Synthetic Data
The recent advances in 3D modeling technology have led to a significant increase in the availability of 3D models with part-level structures accessible online. This progress has opened up new possibilities for detailed analysis and application in various fields, enhancing the resources for researchers and practitioners working with complex 3D structures. Articulated objects are one such class of complex 3D structures that have benefited from this progress. The availability of 3D datasets with part-level structures has enabled the creation of synthetic datasets for articulated objects.
@@
To facilitate the prediction of part mobility via a learning-based approach, hu2017learning collect a synthetic dataset for 3D articulated objects by sourcing shapes from ShapeNet chang2015shapenet and SketchUp sketchup. In this process, they manually segment the shapes into parts and provide detailed annotations for the articulation parameters. Following this foundational work, RPM-Net yan2019rpm extends the scope of the dataset to include a greater variety of objects incorporating more complex motion, e.g., the opening of the umbrella cover. For each shape in the dataset, every possible pair of parts is labeled as a reference part and a moving part. Further, each pair is regarded as a mobility unit, associated with annotated motion parameters. These parameters form a quadruple, including the transformation type (including translation, rotation, or the combination of the two), the position and direction of the transformation axis, and the range of the motion.
@@
Shape2Motion wang2019shape2motion is another synthetic dataset in a larger scale that was introduced concurrently. The shapes are sourced from ShapeNet chang2015shapenet and 3D Warehouse threedwarehouse. Each shape includes part segmentations and articulation parameters that are annotated in a similar way as RPM-Net yan2019rpm and hu2017learning with the use of a developed annotation tool. This tool enhances the efficiency of the annotation process by allowing the annotators to visually verify the correctness by animating the object with the annotated parameters. This allows for boosting the scale of the dataset with a much larger number of objects and movable parts in a wide range of categories.
@@
The datasets mentioned above are valuable resources for research on shape analysis and mobility prediction of articulated objects using 3D meshes or point clouds as input. However, a notable limitation of these datasets is the absence of texture information. This restricts their utility in applications such as simulation, digital twins creation, augmented reality, and other modeling tasks that require a more holistic visual representation. To fill in this gap, the PartNet-Mobility dataset, with a simulation environment SAPIEN  xiang2020sapien, was introduced. This dataset is a subset and an extension from a part-level 3D shape dataset PartNet mo2019partnet, enriched with articulation annotations organized in URDF files. The inclusion of diverse appearances and motions in PartNet-Mobility significantly enhances its value, inspiring increasingly more research in the field of articulated object modeling in connection with visual perception jiang2022opd; sun2023opdmulti; wang2025active and digital twin creation jiang2022ditto; liu2023paris; weng2024neural. Building upon the PartNet-Mobility dataset, OPDSynth jiang2022opd is introduced for the task of openable part detection (OPD) by blending the rendering of the synthetic data with real-world RGB images. This dataset enables the mobility part detection from images, aligning more closely with practical real-world applications.
Later, the GAPartNet geng2023gapartnet dataset is introduced to capture finer-level part details that have been previously overlooked in existing datasets. This dataset is tailored to enhance the generalization of object perception and manipulation tasks across various object categories by focusing on the concept of generalizable and actionable parts. The underlying premise is that functional parts, such as buttons and handles, are fundamental elements whose identification and extraction can significantly improve generalizability within and across object categories. They argue that the functional parts such as buttons and handles are more elementary, and the extraction of these parts can improve the generalizability in an intra-category manner. To achieve a comprehensive and versatile dataset, GAPartNet selectively compiles data from both the PartNet-Mobility xiang2020sapien and AKB-48 liu2022akb datasets to encompass both synthetic and real data. Sourcing from GAPartNet, Drag-a-Move li2024dragapart provides a 2D synthetic dataset for the task of part-oriented drag control. It contains a collection of triplets $(x,y,D)$, where image $x$ and $y$ are the object images in an initial and new states, and $D$ is a collection of drag actions applied to image $x$. The drag annotation is projected from the ground-truth articulation annotations from 3D models to 2D images. This dataset facilitates research on interactive generative models for articulated objects using drag control on 2D images with 3D awareness.
Recently, S2O iliash2024s2o points out that PartNet-Mobility dataset suffers from lack of diversity, contains highly similar objects, and is biased towards objects with simple articulation structures. So they propose a new dataset, Articulated Containers Dataset (ACD), which provides more challenging and realistic 3D objects. The shapes are sourced from ABO collins2022abo, 3D-Future fu20213dfuture, and HSSD khanna2024habitat datasets. The objects in ACD are more complex in geometry (e.g., with L-shaped tables and corner cabinets) and articulation structures with significantly more openable parts. This dataset can potentially help to reduce the gap of models trained on synthetic data generalized to real objects. On the track to increase the data diversity, Arti-PG sun2024artipg introduces a procedural generation toolbox that uses mathematical rules to synthesize articulated objects, with code as the user interface. By randomizing configurations, this toolbox enables unlimited data variations, making it useful for scaling dataset volume.
@@ 3.2 Real-world Data
A primary challenge in working with synthetic data is the unrealistic assumption that it is perfect—overly simplistic, complete, and free from noise. This can result in models that perform well in synthetic environments but struggle when applied to real-world scenarios, where data is often noisy and incomplete To fill this synthetic-real gap, more recent efforts have started to focus on collecting data from real-world sensors, such as RGB-D cameras.
@@
RBO martin2019rbo represents the pioneering dataset in this regard, being the first to collect data from RGB-D scans. It not only provides RGB-D recordings of human interaction with objects under varying experimental conditions, but also creates corresponding synthetic mesh models for each object. ReArt-48 liu2022toward introduced a slightly larger dataset that reconstructs meshes from RGB-D scans and provides the articulation annotations across five object categories. Taking this further, AKB-48 liu2022akb emerges as the first large-scale dataset for articulated objects based on real scans, where each object is described in a knowledge graph. To construct this dataset, a fast articulation knowledge modeling pipeline is presented, significantly reducing the cost and effort required for object modeling in the real world. This innovation enables the creation of 3D models on a scale comparable to PartNet-Mobility. Additionally, AKB-48 also annotates the physical properties of the objects, such as mass, to enhance the dataset’s applicability. This additional annotation is important for bridging the generalization gap between simulation and real-world applications.
@@
In parallel, the MultiScan mao2022multiscan dataset marks a groundbreaking advancement as the first large-scale scene-level dataset that documents multiple states of articulated objects in indoor settings. To compile this dataset, a scalable 3D environment acquisition pipeline is designed for processing raw RGB-D scans to produce 3D surface mesh reconstruction with texture and articulation annotations for each articulated object in the scene. MultiScan is an invaluable resource for advancing research and applications that require an understanding of the dynamics and interactions of articulated objects in real-world environments at the scene level. Leveraging the MultiScan dataset, OPDMulti sun2023opdmulti is introduced for the OPD task of the multiple-parts version by extracting frames from the RGB-D scans and annotating the openable parts in each frame. It extends the OPD task to the scene level, which is more challenging and practical in real-life scenarios.
ParaHome kim2024parahome is a recent real-world dataset that captures 3D human motions and interactions with objects within a home environment. The authors set up a system to capture the human motions with wearable motion capture devices and to track the dynamic interactions with the object synchronized with multi-view RGB cameras. Articulated objects with multiple parts in the dataset are expressed with parameterized articulations. This dataset enables new opportunities for human-object interaction studies, which is an important application in the field of embodied AI and robotics.
@@
To address the task of articulation estimation from 2D, OPDReal jiang2022opd dataset is introduced by collecting RGB-D scans from real-world scenes. By providing a set of real RGB images capturing various articulation states of the objects, along with reconstructed mesh models, OPDReal presents new opportunities for research in the field of articulated object modeling from 2D visuals. OPDMulti sun2023opdmulti follows this trend and extends the task to multiple objects in the scene by collecting RGB-D frames, which is more challenging and practical in real-life scenarios. As a subtask that tackles a segmentation problem for multiple articulated parts, wang2025active introduces a 2D dataset that consists of 2,550 RGB images captured fromthe real world. This dataset focuses on articulated part segmentation by leveraging an active learning model to get fine-grained annotation for part segmentation, saving the cost of manual annotation. However, this dataset does not provide articulation annotations for each part.
@@
geo rep. # states # parts partial/noisy aligned part seg. intermediate rep. strategy supervision geo rep. arti. part seg.
@@
xu2009joint mesh 1 - surface patch handcrafted - mesh
mitra2010illustrating mesh 1 - surface patch handcrafted - mesh
sharf2014mobility mesh multi surface patch handcrafted - mesh
yuan2016space PC multi 3D trajectory handcrafted - PC
li2016mobility RGB-D multi 3D trajectory handcrafted - PC
hu2017learning mesh 1 - surface patch - - mesh
yi2018deep PC 2 PC feat. SL labeled PC PC
Shape2Motion wang2019shape2motion PC 1 - PC feat. SL labeled PC PC
RPM-Net yan2019rpm PC 1 - PC feat. SL labeled PC PC
abbatematteo2019learning RGB-D 1 - image feat. SL prim. params. prim. params.
ANCSH li2020category PC 1 - PC feat. SL labeled PC PC
shi2021self PC multi PC feat. SSL PC PC
MultiBodySync huang2021multibodysync PC multi PC feat. SSL PC PC
CAPTRA weng2021captra PC multi PC feat. SL labeled PC PC
abdul2022learning PC 1 - PC feat. SL labeled PC PC
qian2022understanding RGB multi image feat. SL labeled PC 3D plane
OPD jiang2022opd RGB 1 - - image feat. SL 2D mask 2D mask
OPDMulti sun2023opdmulti RGB 1 - - image feat. SL 2D mask 2D mask
liu2022toward RGB-D 1 - PC feat. SL labeled PC PC
GAPartNet geng2023gapartnet RGB-D 1 - PC feat. SL labeled PC PC
liu2023building PC multi - PC feat. SSL PC PC
liu2023self PC 1 - PC feat. SSL PC PC
liu2023semi PC 1 - PC feat. semi-weakly SL labeled PC PC
Banana deng2023banana PC 1 - PC feat. SL labeled PC PC
wang2025active RGB 1 - image feat. SL 2D mask 2D mask
AutoURDF lin2024autourdf PC multi neural field SSL PC PC
GAMMA yu2024gamma PC 1 - PC feat. SL labeled PC PC
@@
pekelny2008articulated PC multi PC handcrafted - mesh
A-SDF mu2021asdf SDF 1 neural field SL SDF mesh
Ditto jiang2022ditto PC 2 neural field SL labeled PC mesh
CLA-NeRF tseng2022cla MV RGBs 1 neural field SL 2D mask mesh
wei2022self MV RGBs 1 neural field SL MV RGBs mesh
WatchItMove noguchi2022watch MV RGBs multi neural field SSL MV RGBs ellipsoids
CARTO heppert2023carto MV RGBs 1 - neural field SL SDF mesh
liu2023few - - - convex TL - mesh
PARIS liu2023paris MV RGBs 2 neural field SSL MV RGBs mesh
SfA nie2023sfa RGB PC multi neural field SL PC mesh
NAP lei2023nap - - - latent feat. generative - mesh
CAGE liu2024cage - - - latent feat. generative - mesh
weng2024neural MV RGB-Ds 2 neural field SSL MV RGB-Ds mesh
REACTO song2024reacto MV RGBs multi neural field SSL MV RGBs mesh
Real2Code mandi2024real2code RGBs 1 - PC, neural field TL occupancy mesh
RSRD kerr2024robot MV RGBs multi neural field SSL MV RGBs mesh
DragAPart li2024dragapart RGB 1 - image feat. generative RGB RGB
LEIA swaminathan2024leia MV RGBs 4 neural field SSL MV RGBs mesh
PhysPart luo2024physpart - - - latent feat. generative - mesh
SINGAPO liu2024singapo - - - latent feat. generative - mesh
Articulate-Anything le2024articulate - - - latent feat. TL - mesh
MeshArt gao2024meshart - - - latent feat. generative - mesh
ArtFormer su2024artformer - - - latent feat. generative - mesh
Articulate AnyMesh qiu2025anymesh mesh 1 MV RGBs TL - mesh
ArtGS liu2025building MV RGB-Ds 2 neural field SSL MV RGB-Ds mesh
Table 2: Summary of the work in articulated object modeling from the geometric perspective. The table provides information about the geometric representation and the number of articulted states observed in the input, the assumptions made on the input data, the methodology used in the geometry modeling, and the output data representation.
4 Geometry Modeling
In this section, we discuss the recent works on articulated object modeling from the geometry modeling perspective. Based on the goal of the tasks, we categorize the works into two groups: articulated part perception and articulated object creation. The former group of work focuses on understanding the articulated structure of the objects from 3D or 2D observations in terms of the articulated part analysis and articulation parameters estimation, and in this section, we focus the discussion on the subtask of shape analysis in each work. The latter group focuses on reconstructing or generating the 3D shape of objects either with motion parameters associated for each part or with the object animated into different articulation states over time. In Table 2, we summarize these works in terms of the choice of geometric representation, the assumptions that each work made for the input, and the methodology each work adopted for the articulated shape analysis. We specify the meaning of each column in Table 2 as follows: